# SL2A-INR: Single-Layer Learnable Activation for Implicit Neural Representation

### v3

@@ 6.3 Robustness to Hyperparameter Settings
6.4 Activation Function Visualization
@@ ACKNOWLEDGMENTS
This supplementary material presents further details, including the specific hyperparameter values, the reasoning behind the proposed method, an analysis of the computational complexity of SL2A, and additional applications of our approach to neural radiance fields and single image super-resolution. These examples underscore the robustness of our method and demonstrate its adaptability across various domains. Additionally, we provide the code for our paper, along with baseline implementations, in an anonymous format to enhance the reproducibility of our results.
@@ 8 SL2A Design Choice
Method $N_{\theta}$ Time (min.) Size (MB) PSNR SSIM
@@
As shown in Fig. 9, removing ReLU results in a consistent PSNR drop, with more severe degradation observed in lower-degree Chebyshev expansions, highlighting the critical role of ReLU in enhancing the model’s expressive power to capture complex signal representations. The largest performance drop is 6.22 dB in lower-degree configurations, demonstrating its critical importance, while even in higher-degree settings, ReLU provides measurable improvements, reinforcing its necessity for expressivity. These results confirm that the modulation of layers by learnable activation $\Psi(\mathbf{x})$ could be insufficient and introducing nonlinearity such as ReLU improves performance and expressive power.
@@ 9 Chebyshev Degree/ Linear layer rank Efeect
Figures 10 and 11 illustrate the impact of rank of linear layers and the polynomial degree on PSNR for image representation tasks. As observed, increasing both the Chebyshev polynomial degree and rank of linear layers leads to higher PSNR values, demonstrating the effectiveness of higher-degree representations in capturing finer details. More importantly, the results reveal a critical trade-off: when rank and model size are constrained, increasing the polynomial degree in the initial layer leads to better accuracy compared to configurations with higher rank but a lower degree. For instance, the blue point on the right (363k parameters) achieves superior PSNR than the orange point in the middle (330k parameters) despite having fewer rank, and similar trends hold across other configurations. This validates our architectural choice of employing lower-rank linear layers in deeper layerswhile allowing a higher polynomial degree in the initial layer, enabling efficient representations without excessive parameter growth. Figure 11 further supports this intuition by presenting the same analysis without ReLU activations. The observed trends remain consistent, reinforcing that altering the rank of linear layers alone is less effective in improving performance when ReLU is removed. This highlights the complementary role of activation functions and structured polynomial expansions in achieving a balance between efficiency and expressivity in our proposed architecture.
@@
10 Neural Radiance Fields
Neural Radiance Fields (NeRFs) [23] provide a method for synthesizing novel views of complex 3D scenes by representing them as continuous volumetric functions. These functions are optimized using 2D images and corresponding camera poses. The NeRF architecture employs a Multi-Layer Perceptron (MLP) to implicitly encode the scene. It takes 5D inputs—comprising a 3D spatial location and a 2D viewing direction—and outputs the associated RGB color and volume density. For training, we adopt the same configurations as FINER [19] for all baselines and report their corresponding results accordingly. Table 5 presents a quantitative comparison of our method against various methods on the Blender dataset [23]. SL2A outperforms competing approaches in nearly all scenes and remains highly competitive in the few cases where it does not achieve the best results. This demonstrates SL2A’s ability to deliver consistently strong performance across diverse tasks.
Table 5: Quantitative comparisons on novel view synthesis.
Methods Chair Drums Ficus Hotdog Lego Materials Mic Ship
Gauss 32.68 23.16 26.10 32.17 28.29 26.19 33.59 22.28
SIREN 33.31 24.89 27.26 32.85 29.60 27.13 33.28 22.25
WIRE 29.31 22.22 25.91 30.11 25.76 25.05 32.35 21.15
FINER 33.90 24.90 28.70 33.05 30.04 27.05 33.96 22.47
SL2A 35.23 25.07 28.33 34.32 31.16 29.35 34.75 23.16
@@ 11 Hyperparameter Settings
Following prior studies configuration, we utilized the hyperparameter settings shown in  Tab. 6 to ensure consistency. Additionally, we adopted their initialization scheme, using the code publicly released by the authors.
@@ 12 Computational Complexity
We analyze the computational complexity in terms of the number of parameters, training time (per image/SDF), and model size (MB) across different tasks, as summarized in Tabs. 7 and 9. Although our method, SL2A, utilizes more parameters compared to FINER (the best-performing method after ours), it does not require extensive training time or significant associated costs. Our primary contribution is the elimination of the need for manual activation function design, in contrast to prior approaches. Moreover, as illustrated in Tab. 4, existing methods utilizing naive learnable activations, such as KAN [18], suffer from limited scalability and demand substantial training resources. We effectively address these limitations by introducing a scalable learnable activation paradigm that requires only marginally increased parameter counts. Future research will explore further enhancements aimed at improving network compactness. In the image representation task (Tab. 7), SL2A achieves competitive training time, considerably outperforming many existing methods, and remains efficient despite its marginally higher parameter count. For the occupancy representation task (Tab. 9), SL2A requires slightly longer training times compared to FINER but still maintains efficiency within acceptable bounds. Overall, the complexity analysis indicates that SL2A remains computationally efficient across tasks, with slight trade-offs in training time balanced by its superior performance in representation quality. We observed similar computational trends in the neural field representation task as well.
@@
Table 7: Comparison of computational complexity across different methods (image representation task).
Method #Params (K)$\downarrow$ Training Time (min.) $\downarrow$ Size (MB) $\downarrow$
FINER 0.198 0.595 0.80
WIRE 0.099 1.713 0.90
Gauss 0.198 3.08 1.50
SIREN 0.198 0.643 1.20
ReLU+P.E. 0.204 3.425 1.00
SL2A 0.330 0.77 0.93
@@
Table 8: Computational complexity for Occupancy Representation Task.
Method #Params (K)$\downarrow$ Training Time (min.) $\downarrow$ Size (MB) $\downarrow$
FINER 0.198 42.41 0.80
WIRE 0.066 63.03 0.53
Gauss 0.198 32.05 0.80
SIREN 0.198 27.22 0.80
ReLU+P.E. 0.214 39.32 0.86
SL2A 0.248 45.59 1.00
Table 9: Computational complexity comparison for Occupancy Representation tasks.
13 Inverse Problems
@@
To demonstrate the generalization capability of our approach to inverse problems, we evaluated it on the task of single image super-resolution using an image from the DIV2K dataset [40]. Specifically, we trained our architecture for image representation and downsampled the original image (with dimensions $1356\times 2040\times 3$) by factors of $2$, $4$, and $6$ and evaluated it on the original resolution. The corresponding results, illustrated in Fig. 12, compare super-resolution reconstructions of a parrot image across multiple ratios. Our findings indicate that SL2A-INR consistently surpasses FINER, achieving superior PSNR [15] metrics. Furthermore, SL2A-INR distinctly preserves sharp details and produces less noisy result images, specifically in 6× settings, whereas the recent SOTA method FINER produces comparatively noisier results. These visual and quantitative improvements highlight our method’s effectiveness in achieving high-quality super-resolution reconstructions and underscore its robustness and adaptability to inverse problems. We hypothesize that this trend could similarly hold for other inverse problem tasks involving complex reconstructions, a direction we plan to explore in future work.

### v4

@@ 6.3 Robustness to Hyperparameter Settings
7 Limitations and Future Work
SL2A builds on prior work (KAN) [17], improving its efficiency by using Chebyshev polynomials in the first layer and incorporating a hybrid architecture with low-rank linear layers. However, scaling it to very large architectures, as inherited from KANs, can be costly. Moreover, in the image-fitting task, SL2A introduces additional, though small, parameters. Incorporating sparsification or weight sharing represents a promising future direction to address these challenges.
@@ ACKNOWLEDGMENTS
This supplementary material presents further details, including the activation function visualization, the specific hyperparameter values, the reasoning behind the proposed method, an analysis of the computational complexity of SL2A, and additional applications of our approach to single image super-resolution.
9 Activation Function Visualization
@@ 10 SL2A Design Choice
Method #Params (M) Time (min.) Size (MB) PSNR SSIM
@@
$f_{\theta}(\mathbf{x})$ $=\mathbf{W}_{L}(\mathbf{z}_{L-1}\odot\mathbf{z}_{1})+\mathbf{b}_{L}.$
As shown in Fig. 8, removing ReLU results in a consistent PSNR drop, with more severe degradation observed in lower-degree Chebyshev expansions, highlighting the critical role of ReLU in enhancing the model’s expressive power to capture complex signal representations. The largest performance drop is 6.22 dB in lower-degree configurations, demonstrating its critical importance, Even in higher-degree settings, ReLU continues to provide improvements, reinforcing its necessity for expressivity. These results confirm that the modulation of layers by learnable activation $\Psi(\mathbf{x})$ could be insufficient and introducing nonlinearity such as ReLU improves performance and expressive power.
@@ 11 Chebyshev Degree/ Linear layer rank Efeect
Figures 9 and 10 illustrate the impact of rank of linear layers and the polynomial degree on PSNR for image representation tasks. As shown Fig. 9, increasing both the Chebyshev polynomial degree and rank of linear layers leads to higher PSNR values, demonstrating the effectiveness of higher-degree representations in capturing finer details. More importantly, the results reveal a critical trade-off: under similar model size or parameter constraints, increasing the polynomial degree in the initial layer generally yields better accuracy than using higher-rank configurations with a lower degree in the image representation task. For instance, the blue point on the right (363k parameters) achieves superior PSNR than the orange point in the middle (330k parameters) despite having fewer rank, and similar trends hold across other configurations. This validates our architectural choice of employing lower-rank linear layers in deeper layers while allowing a higher polynomial degree in the initial layer, enabling efficient representations without excessive parameter growth. Figure 10 further supports this intuition by presenting the same analysis without ReLU activations. The observed trends remain consistent, reinforcing that altering the rank of linear layers alone is less effective in improving performance when ReLU is removed. This highlights the complementary role of non-linear activation functions and low-rank layers in achieving a balance between efficiency and expressivity in our proposed architecture.
@@ 12 Hyperparameter Settings
In accordance with the configurations used in prior studies, we employed the hyperparameter settings summarized in Tab. 5 for the image fitting task. Furthermore, we adopted the same initialization scheme as in these works, utilizing the publicly available implementation provided by the authors.
@@ 13 Computational Complexity
We analyze the computational complexity in terms of the number of parameters, and training time (per image/SDF), across different tasks, as summarized in Tabs. 6 and 7. Although our method, SL2A, utilizes more parameters compared to FINER (the best-performing method after ours), it does not require extensive training time. It is important to note that the reported training times correspond to a fixed number of iterations using optimal hyperparameters that maximize performance (PSNR) rather than configurations optimized for speed. Time differences between methods arise from their respective optimal hyperparameter settings. Our primary contribution is the elimination of the need for manual activation function design, in contrast to prior approaches. Moreover, as illustrated in Tab. 4, existing methods utilizing naive learnable activations, such as KAN [17], suffer from limited scalability and demand substantial training resources. We effectively address these limitations by introducing a scalable learnable activation paradigm that requires only marginally increased parameter counts while maintaining computational efficiency through the use of a single learnable activation layer and low-rank MLPs, which minimize overhead without sacrificing performance. Future research will explore further enhancements aimed at improving network compactness. In the image representation task (Tab. 6), SL2A achieves competitive training time, considerably outperforming many existing methods, and remains efficient despite its marginally higher parameter count. For the occupancy representation task (Tab. 7), SL2A requires slightly longer training times compared to FINER but still maintains efficiency within acceptable bounds. Overall, the complexity analysis indicates that SL2A remains computationally efficient across tasks, with slight trade-offs in training time balanced by its superior performance in representation quality.
@@
Table 6: Comparison of computational complexity across different methods for the Image Representation Task.
[…]
FINER 0.198 0.595
WIRE 0.099 1.713
Gauss 0.198 3.08
SIREN 0.198 0.643
ReLU+P.E. 0.204 3.425
SL2A 0.330 0.77
@@
Table 7: Computational complexity across different methods for the Occupancy Representation Task.
[…]
FINER 0.198 42.41
WIRE 0.066 63.03
Gauss 0.198 32.05
SIREN 0.198 27.22
ReLU+P.E. 0.214 39.32
SL2A 0.248 45.59
14 Single Image Super Resolution
@@
To demonstrate the generalization capability of our approach, we evaluated it on the task of single image super-resolution using an image from the DIV2K dataset [39]. Specifically, we trained our architecture for image representation and downsampled the original image (with dimensions $1356\times 2040\times 3$) by factors of $2$, 44, and $6$ and evaluated it on the original resolution. The corresponding results, illustrated in Fig. 11, compare super-resolution reconstructions of a parrot image across multiple ratios. Our findings indicate that SL2A-INR consistently surpasses FINER, achieving superior PSNR [14] metrics. Furthermore, SL2A-INR distinctly preserves sharp details and produces less noisy result images, specifically in 6× settings, whereas the recent SOTA method FINER produces comparatively noisier results. These visual and quantitative improvements highlight our method’s effectiveness in achieving high-quality super-resolution reconstructions and underscore its robustness and adaptability to inverse problems. We hypothesize that this trend could similarly hold for other inverse problem tasks involving complex reconstructions, a direction we plan to explore in future work.